Consider the following data from a repeated-measur
Answer #1

Answer

The completed table is as follows

Observation A B D=B-A D^2
1 1 2 1 1
2 3 3 0 0
3 4 6 2 4
4 6 5 -1 1
5 6 8 2 4
Total 4 10

Hence the mean difference score is M_D=4/5=0.8

Now estimated standard deviation of the different scores is

s=\sqrt{\frac{\sum D^2-n.M_D^2}{n-1}}=\sqrt{\frac{6.8}{4}}=1.30384

Now the estimated standard error of the mean difference scores,

s_{MD}=\frac{s}{\sqrt{n}}=0.5831

The t statistic to test the null hypothesis H_0:\mu_D=0

t=\frac{M_D-\mu_0}{s_{MD}} here \mu_0 is the value of \mu_D under Ho , hence 0

  =\frac{0.8}{0.5831}=1.37

Now to conduct a two tailed test the degrees of freedom are n-1=5-1=4

At level of significance a-0.05 the critical values are -2.78 and 2.78

Finally, since the t-statistic does not belong to the critical region (since -2.78<1.37<2.78) Hence we accept null hypothesis based on the data.

Hnce the answer...

Thank you......